# Do Language Models Understand the Cognitive Tasks Given to Them? Investigations with the N-Back Paradigm

### v2

@@ 4.4 T1 Model Performance as N Increases
Given that the best model, Llama-3.1-70b-Instruct, performs well for 1 through 3-back tasks, we would like to know how its performance might change for larger n’s.
Figure 9 shows that the retrieval accuracy gradually declines as n increases; although, even at $n=8,9,10,$ the model is still able to exactly retrieve the correct letters 75.25%, 66.08%, and 57.1% of the times, which translates to task accuracies of 83.33%, 78.25%, and 71.92%.
In addition, we measure $\mathsf{P}_{n,m}$ for each $n,m\in\{1,2,3,...,10\}$, as shown in Figure 10. We notice that $\max_{m}\mathsf{P}_{n,m}=\mathsf{P}_{n,n}$ for $101\leq n<$ 10. Moreover, $\mathsf{P}_{n,m}$ tends to decrease symmetrically as $m$ deviates from $n$. We can consider this pattern as a signature of n-back task understanding.
@@ 4.5 Curriculum Learning
The practice of training models on examples of increasing difficulty is known in machine learning as curriculum learning (Bengio et al., 2009). Here, we repeat the experiments from Section 4.4 with in-context curriculum learning to gradually familiarize the model with the task. Specifically, before prompting Llama 3.1 70b Instruct to perform an n-back task, we provide demonstrations that include letter sequences and corresponding correct responses for tasks ranging from 1-back to n-back. As shown in Figure 9, this approach leads to significant improvements in performance for larger n values. The model achieves retrieval accuracies of 79.83%, 80.17%, and 71.67% and task accuracies of 90.08%, 90.08%, and 84.75%, for $n=8,9,10$.
@@
Table 2: Retrieval accuracies on 2-back and 3-back tasks, for representative models, under interactive demo.
Figure 11: Retrieval accuracies under alternative answer format on 1-back, 2-back, and 3-back tasks, averaged across 50 trials, for all models tested, organized by model family.
@@
4.6 Interactive Demo
For both 2-back and 3-back trials, Qwen 1.5 14b Chat fails to achieve two consecutive correct answers after 10 demo sequences, further confirming its difficulty with task comprehension. Interestingly, Gemma 2 27b Instruct performs better in 2-back trials compared to the original experiments (non-interactive demo) but does worse on 3-back trials, as shown in Table 2. For Llama 3.1 70b Instruct, 2-back performance remains high at 99% with the interactive demo but drops significantly for 3-back trials.
4.7 Reciting N Most Recent Stimuli
We experiment with an alternative answer format that encodes task requirements in greater detail. For 2-back trials, models are instructed to answer “current: [ current letter ], 1 back: [ letter 1 back ], 2 back: [ letter 2 back ]; current letter [ current letter ] and letter 2 back [ letter 2 back ] are [ different / identical ].” For 3-back trials, they are instructed to answer “current: [ current letter ], 1 back: [ letter 1 back ], 2 back: [ letter 2 back ], 3 back: [ letter 3 back ]; current letter [ current letter ] and letter 3 back [ letter 3 back ] are [ different / identical ].” Retrieval accuracies show significant improvements (Figure 11), including for T3 models, though their performance still lags slightly on 3-back trials.
We include these results only for comparison, given that this format changes the original task into one that allows covert verbal rehearsal. In human experiments, participants would not have enough time to recite all n most recent letters upon presentation of each new letter. However, these results do highlight the malleability of language models’ performance on working memory tasks.
@@ 4.8 Attention Analysis
Attentions in transformer-based language models reveal how much each generated token attends to every preceding token. We hypothesize that, for each retrieval, a more performant model should attend more to the source token from n steps back. This is precisely what we observe in the Qwen models. For each (trial, layer, head), we obtain the mean retrieval attention (MRAT) by averaging the attention each retrieval token gives to the correct source token. Indeed, compared to the 14b model, Qwen 2 72b Instruct (T1) shows a higher concentration at the tail end of the MRAT distribution (Figure 12). Figure 13 shows the attention map with the highest MRAT, which almost exactly matches our hypothesized pattern. However, Llama models do not exhibit this pattern to the same degree. Attentions in Llama 3.1 models are much more diffuse. The maximum MRATs for Llama 3.1 8b Instruct and Llama 3.1 70b Instruct are 4.86% and 8.52%, respectively.
@@ 5 Conclusion
In this work, we apply the n-back task, a common working memory test, to a range of language models. We categorize the models into three performance tiers based on retrieval accuracy. Through analyzing a representative model from each tier, we identify three distinct levels of task understanding and task set maintenance capabilities, highlighting their confounding nature. We challenge the best model to perform 1 through 10-back tasks, noticing a signature of task comprehension and the benefit of in-context curriculum learning for larger n’s. We find that interactive demos, though closer to human study paradigms, are less effective at conveying the task. For comparison, we also include results under an alternative answer format that allows for rehearsal. Finally, we observe that more performant models tend to have higher retrieval attentions, although the values for Llama models seem surprisingly small.

### v3

@@ 4.4 T1 Model Performance as N Increases
Given that the best model, Llama-3.1-70b-Instruct, performs well for 1 through 3-back tasks, we would like to know how its performance might change for larger $n$’s. Figure 9 shows that the retrieval accuracy gradually declines as $n$ increases; although, even at $n=8,9,10,$ the model is still able to exactly retrieve the correct letters 75.25%, 66.08%, and 57.1% of the time, which translates to task accuracies of 83.33%, 78.25%, and 71.92%. In addition, we measure $\mathsf{P}_{n,m}$ for each $n,m\in\{1,2,3,...,10\}$, as shown in Figure 10. We notice that $\max_{m}\mathsf{P}_{n,m}=\mathsf{P}_{n,n}$ for $101\leq n<$ 10. Moreover, $\mathsf{P}_{n,m}$ tends to decrease symmetrically as $m$ deviates from $n$. We argue that this pattern points to true $n$-back task understanding.
@@ 4.5 Curriculum Learning
The practice of training models on examples of increasing difficulty is known in machine learning as curriculum learning (Bengio et al., 2009). Here, we repeat the experiments from Section 4.4 with in-context curriculum learning to gradually familiarize the model with the task. Specifically, before prompting Llama 3.1 70b Instruct to perform an $n$-back task, we provide instructions and demonstrations that include letter sequences and corresponding correct responses for tasks ranging from 1-back to $n$-back. For example, to prepare the model for the 4-back task, we prepend 4 complete demonstration sequences (1-back to 4-back) to the context, before starting the test sequence. As shown in Figure 9, this approach leads to significant improvements in performance for larger $n$ values. The model achieves retrieval accuracies of 79.83%, 80.17%, and 71.67% and task accuracies of 90.08%, 90.08%, and 84.75% for $n=8,9,10$.
@@
Table 2: Retrieval accuracies on 2-back and 3-back tasks, for representative models, with interactive demos.
4.6 Interactive Demo
For both 2-back and 3-back tasks, Qwen 1.5 14b Chat (T3) fails to achieve two consecutive correct answers after 10 demo sequences, further confirming the model’s difficulty with task comprehension. The complete 2-back dialogue is included in Appendix A. Interestingly, Gemma 2 27b Instruct (T2) performs better on 2-back trials compared to the original experiments but does worse on 3-back trials, as shown in Table 2. Llama 3.1 70b Instruct (T1) maintains high performance at 99% on 2-back trials but shows a significant drop in performance on 3-back.
@@ 4.7 Attention Analysis
Attentions in transformer-based language models reveal how much each generated token attends to every preceding token. We hypothesize that, for each retrieval, a more performant model should attend more to the source token from $n$ steps back. This is precisely what we observe in the Qwen models. For each (trial, layer, head), we obtain the mean retrieval attention (MRAT) by averaging the attention each retrieved token gives to the correct source token. Compared to the 14b model, Qwen 2 72b Instruct (T1) contains a much larger proportion of high-MRAT attentions (Figure 11), with its highest scoring attention (71.98%) closely matching our hypothesized pattern (Figure 12). However, Llama models do not exhibit this pattern to the same degree. Attentions in Llama 3.1 models are much more diffuse. The maximum MRATs for Llama 3.1 8b Instruct and Llama 3.1 70b Instruct are 4.86% and 8.52%, respectively.
@@ 5 Conclusion
In this work, we apply the $n$-back task, a common working memory test, to a range of language models, identifying three distinct performance tiers. We find that these tiers differ not only in retrieval accuracy but also in our measure of task understanding and task set maintenance, suggesting that the performance gap is due at least in part to these differences. We challenge the best model to perform 1 through 10-back tasks, noticing a signature of task comprehension and the benefit of in-context curriculum learning for larger $n$’s. We find that interactive demos, though closer to human study paradigms, are less effective at conveying the task. Finally, we notice that more performant models tend to have higher retrieval attentions.
@@ 6 Limitations
7 Acknowledgments
We thank the anonymous reviewers for their helpful feedback. XH would also like to acknowledge the participants in the Brown MICAAD study, who partially inspired this project.
Appendix A Interactive Demo Dialogue
Qwen 1.5 14b Chat (T3) 2-back interactive demo dialogue.
<|im_start|>system
You are a participant in a cognitive task. You will
be shown a sequence of letters, one at a time. For
each letter, determine if it is the same as the
letter 2 steps before it. Answer in the following
format: ’[current letter] and [letter 2 back] are
[identical/different].’ Note that, for the first 2
letters, there won’t be any letter 2 steps back;
write ’none’ for ’[letter 2 back]’ in this case.
Let’s think step by step.
<|im_end|>
<|im_start|>user
For example, given the sequence t, z, h, z,
the answers should be:
t and none are different.
z and none are different.
h and t are different.
z and z are identical.
Now, given the sequence t, z, t, h,
what should the answers be?
<|im_end|>
<|im_start|>assistant
t and none are different.
z and t are different.
t and z are different.
h and t are different.
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
t and none are different.
(There was no letter 2 steps ago.)
z and none are different.
(There was no letter 2 steps ago.)
t and t are identical.
(The letter 2 steps ago was t.)
h and z are different.
(The letter 2 steps ago was z.)
Now, given the sequence v, n, c, n,
what should the answers be?
<|im_end|>
<|im_start|>assistant
v and none are different.
n and v are different.
c and n are different.
n and n are identical.
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
v and none are different.
(There was no letter 2 steps ago.)
n and none are different.
(There was no letter 2 steps ago.)
c and v are different.
(The letter 2 steps ago was v.)
n and n are identical.
(The letter 2 steps ago was n.)
Now, given the sequence v, n, v, c,
what should the answers be?
<|im_end|>
<|im_start|>assistant
v and none are different.
n and v are different.
v and n are different.
c and v are different.
(The letter 2 steps ago was v.)
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
v and none are different.
(There was no letter 2 steps ago.)
n and none are different.
(There was no letter 2 steps ago.)
v and v are identical.
(The letter 2 steps ago was v.)
c and n are different.
(The letter 2 steps ago was n.)
Now, given the sequence k, l, b, l,
what should the answers be?
<|im_end|>
<|im_start|>assistant
k and none are different.
l and k are different.
b and l are different.
l and l are identical.
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
k and none are different.
(There was no letter 2 steps ago.)
l and none are different.
(There was no letter 2 steps ago.)
b and k are different.
(The letter 2 steps ago was k.)
l and l are identical.
(The letter 2 steps ago was l.)
Now, given the sequence k, l, k, b,
what should the answers be?
<|im_end|>
<|im_start|>assistant
k and none are different.
l and k are different.
k and l are different.
b and k are different.
(The letter 2 steps ago was k.)
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
k and none are different.
(There was no letter 2 steps ago.)
l and none are different.
(There was no letter 2 steps ago.)
k and k are identical.
(The letter 2 steps ago was k.)
b and l are different.
(The letter 2 steps ago was l.)
Now, given the sequence r, f, m, f,
what should the answers be?
<|im_end|>
<|im_start|>assistant
r and none are different.
f and r are different.
m and f are different.
f and f are identical.
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
r and none are different.
(There was no letter 2 steps ago.)
f and none are different.
(There was no letter 2 steps ago.)
m and r are different.
(The letter 2 steps ago was r.)
f and f are identical.
(The letter 2 steps ago was f.)
Now, given the sequence r, f, r, m,
what should the answers be?
<|im_end|>
<|im_start|>assistant
r and none are different.
f and r are different.
r and f are different.
m and r are different.
(The letter 2 steps ago was r.)
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
r and none are different.
(There was no letter 2 steps ago.)
f and none are different.
(There was no letter 2 steps ago.)
r and r are identical.
(The letter 2 steps ago was r.)
m and f are different.
(The letter 2 steps ago was f.)
Now, given the sequence y, p, w, p,
what should the answers be?
<|im_end|>
<|im_start|>assistant
y and none are different.
p and y are different.
w and p are different.
p and p are identical.
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
y and none are different.
(There was no letter 2 steps ago.)
p and none are different.
(There was no letter 2 steps ago.)
w and y are different.
(The letter 2 steps ago was y.)
p and p are identical.
(The letter 2 steps ago was p.)
Now, given the sequence y, p, y, w,
what should the answers be?
<|im_end|>
<|im_start|>assistant
y and none are different.
p and y are different.
y and p are different.
w and y are different.
(The letter 2 steps ago was y.)
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
y and none are different.
(There was no letter 2 steps ago.)
p and none are different.
(There was no letter 2 steps ago.)
y and y are identical.
(The letter 2 steps ago was y.)
w and p are different.
(The letter 2 steps ago was p.)
Now, given the sequence s, j, g, j,
what should the answers be?
<|im_end|>
<|im_start|>assistant
s and none are different.
j and s are different.
g and j are different.
j and j are identical.
<|im_end|>
<|im_start|>user
This is incorrect.
The answers should be:
s and none are different.
(There was no letter 2 steps ago.)
j and none are different.
(There was no letter 2 steps ago.)
g and s are different.
(The letter 2 steps ago was s.)
j and j are identical.
(The letter 2 steps ago was j.)
Now, given the sequence s, j, s, g,
what should the answers be?
<|im_end|>
<|im_start|>assistant
s and none are different.
j and s are different.
s and j are different.
g and s are different.
(The letter 2 steps ago was s.)
<|im_end|>
Figure 13: Average retrieval accuracies on 1-, 2-, and 3-back tasks, grouped by model family.
Appendix B Reciting N Most Recent Stimuli
We experiment with an alternative answer format that encodes task requirements in greater detail. For 2-back trials, models are instructed to answer “current: {current letter}, 1 back: {letter 1 back}, 2 back: {letter 2 back}; current letter {current letter} and letter 2 back {letter 2 back} are {different / identical}.” The 3-back format is analogous.
Retrieval accuracies show significant improvements, including for T3 models, though their performances still lag slightly on 3-back trials (Figure 13). We include these results only for comparison, given that this format changes the original task into one that allows covert verbal rehearsal. In human experiments, participants would not have enough time to recite all $n$ most recent letters upon presentation of each new letter. However, these results do highlight the malleability of language models’ performance on working memory tasks.